A 40 kg crate is pulled 50 m along a horizontal floor by a constant force exerted by a person, Fp = 150 N, which acts as 30 degree angle. The floor is rough and exerts a friction for Ff = 40 N. Determine: a) the work done by each force acting on the crate b) the net work done on the crate
Answer #1

Work done by the force acting on the crate

Wp = Fp.s = Fp*s*cos theta

Wp = 150*50*cos 30 deg = 6495.19 J

B.

work due to normal force

Wn = Fn*s*cos theta

theta = 90 deg

cos 90 = 0

Wn = 0

Wg = Fg*s*cos theta

theta = 90 deg

cos 90 = 0

Wg = 0 J

total work will be

Wt = Wp + Wn + Wg

Wt = 6495.19 J